# MetaML-Pro: Cross-Stage Design Flow Automation for Efficient Deep Learning Acceleration

### v2

@@ 5.8. Tolerable Loss Variation
Table 4. Performance comparison with the FPGA designs of Jet-DNN network using other approaches on Xilinx FPGAs with a clock frequency of 200MHz.
Model $\alpha_{s}$, $\alpha_{p}$, $\alpha_{q}$ (%) FPGA Acc.(%) Lat. (ns) DSP (%) LUT (%)
HLS4ML Jet-DNN (Duarte et al., 2018) - KU115 75 75 954 (17.3) -
FPL’20 LogicNets JSC-M (Umuroglu et al., 2020) - VU9P 70.6 NA 0 (0) 14,428 (1.2)
FPL’20 LogicNets JSC-L (Umuroglu et al., 2020) - VU9P 71.8 13a 0 (0) 37,931 (3.2)
NMI’21 (Coelho et al., 2021) Qkeras Q6 - VU9P 74.8 55 124 (1.8) 39,782 (3.4)
NMI’21 (Coelho et al., 2021) AutoQkeras QE - VU9P 72.3 55 66 (1.0) 9,149 (0.8)
NMI’21 (Coelho et al., 2021) AutoQkeras QB - VU9P 71.9 70 69 (1.0) 11,193 (0.9)
PL’23 (Que et al., 2023) MetaML - VU9P 76.1 70 638 (9.3) 69,751 (5.9)
FPL’23 (Que et al., 2023) MetaML {0.05, 2, 1} VU9P 75.6 45 50 (0.7) 6,698 (0.6)
FPL’23 (Que et al., 2023) MetaML {0.05, 2, 4} VU9P 72.8 40 23 (0.2) 7,224 (0.6)
This work (Best Acc.) {0.5, 0.1, 0.1} VU9P 76.2 55 272 (4.0) 14,580 (1.2)
This work (Best DSP) {0.5, 3, 4} VU9P 69.7 40 5 (0.1) 9,026 (0.8)
This work (Best LUT) {2, 5, 1} VU9P 72.1 45 47 (0.7) 4,410 (0.4)
This workb (Acc.-DSP-LUT) {0.5, 2, 0.5} VU9P 76.1 50 70 (1.0) 13,042 (1.1)
This workb (Acc.-DSP-LUT) {4, 0.5, 0.5} VU9P 75.7 45 50 (0.7) 6,634 (0.6)
This workc (Acc.-LUT) {4, 4, 0.5} VU9P 73.7 45 54 (0.8) 5,624 (0.5)
This workd (Acc.-DSP) {2, 3, 4} VU9P 70.9 40 12 (0.2) 7,637 (0.8)
@@
b On both Accuracy-DSP and Accuracy-LUT Pareto lines as shown in Fig 20.
@@ 5.9. Design-Space Exploration Strategies
Finding the optimal designs with grid search is time-consuming, as discussed in the previous section. This section investigates various DSE algorithms, including grid search, stochastic grid search (SGS), and Bayesian optimization. Fig. 19 presents the results. Each colored dot represents an design at a specific iteration, with different colors indicating different algorithms. The solid lines represent the Pareto Frontier of each algorithm over iterations, with each colored line corresponding to 22 iterations taking 40 hours. The uppermost grey line indicates a total of 343 iterations using extensive grid search over 624 hours, representing a baseline for comparison. The Bayesian optimization achieves similar results with just 22 iterations, significantly reducing processing time by a factor of 15.6 compared to Grid Search. Compared to SGS, Bayesian optimization’s efficient parameter search yields multiple points near the Pareto frontier, indicating its effectiveness in finding optimal designs and approaching global optima. This effective parameter search by employing past results demonstrates the robustness of the Bayesian optimization approach, enabling users to optimize models while minimizing time and effort.
5.10. Discussion and Comparison
@@ 6. Conclusion
This paper presents a novel co-optimization framework for FPGA-based DNN accelerators, which comprises building blocks that facilitate rapid development of customized cross-stage design flows, automating the entire design iteration process. The results demonstrate that our approach significantly reduces DSP resource usage by up to 92% and LUT usage by up to 89%, while maintaining accuracy and without requiring human effort or domain expertise. In addition, our results reveal that Bayesian optimization in DSE significantly streamlines the process, achieving results comparable to grid search but with a 15.6-fold reduction in processing time. Future work will extend this approach to include RTL space, support more DNN architectures, including Variational Autoencoder (VAE) (Que et al., 2024b) and transformer (Wojcicki et al., 2022) networks, and incorporate more optimization strategies like balancing initiation interval (Que et al., 2021; Rognlien et al., 2022) for further improvements in hardware efficiency and model performance.
Acknowledgement: The support of the United Kingdom EPSRC (grant number UKRI256, EP/V028251/1, EP/N031768/1, EP/S030069/1, and EP/X036006/1), Intel, and AMD is gratefully acknowledged. We thank Markus Rognlien, Shuo Liu and Anyan Zhao for their help.

### v3

@@ 5.8. Tolerable Loss Variation
Table 4. Performance comparison with the FPGA designs of Jet-DNN network using other approaches on Xilinx FPGAs with a clock frequency of 200MHz. Please note that this table only reports Jet-DNN to enable an apples-to-apples comparison with prior Jet-DNN FPGA designs.
Model $\alpha_{s}$, $\alpha_{p}$, $\alpha_{q}$ (%) FPGA Acc.(%) Lat. (ns) DSP (%) LUT (%) Pow. (W)e
HLS4ML Jet-DNN (Duarte et al., 2018) - KU115 75 75 954 (17.3) - -
FPL’20 LogicNets JSC-M (Umuroglu et al., 2020) - VU9P 70.6 NA 0 (0) 14,428 (1.2) -
FPL’20 LogicNets JSC-L (Umuroglu et al., 2020) - VU9P 71.8 13a 0 (0) 37,931 (3.2) -
NMI’21 (Coelho et al., 2021) Qkeras Q6 - VU9P 74.8 55 124 (1.8) 39,782 (3.4) -
NMI’21 (Coelho et al., 2021) AutoQkeras QE - VU9P 72.3 55 66 (1.0) 9,149 (0.8) -
NMI’21 (Coelho et al., 2021) AutoQkeras QB - VU9P 71.9 70 69 (1.0) 11,193 (0.9) -
FPL’23 (Que et al., 2023) MetaML - VU9P 76.1 70 638 (9.3) 69,751 (5.9) 2.506
FPL’23 (Que et al., 2023) MetaML {0.05, 2, 1} VU9P 75.6 45 50 (0.7) 6,698 (0.6) 0.199
FPL’23 (Que et al., 2023) MetaML {0.05, 2, 4} VU9P 72.8 40 23 (0.2) 7,224 (0.6) 0.169
This work (Best Acc.) {0.5, 0.1, 0.1} VU9P 76.2 55 272 (4.0) 14,580 (1.2) 1.096
This work (Best DSP) {0.5, 3, 4} VU9P 69.7 40 5 (0.1) 9,026 (0.8) 0.118
This work (Best LUT) {2, 5, 1} VU9P 72.1 45 47 (0.7) 4,410 (0.4) 0.204
This workb (Acc.-DSP-LUT) {0.5, 2, 0.5} VU9P 76.1 50 70 (1.0) 13,042 (1.1) 0.386
This workb (Acc.-DSP-LUT) {4, 0.5, 0.5} VU9P 75.7 45 50 (0.7) 6,634 (0.6) 0.247
This workc (Acc.-LUT) {4, 4, 0.5} VU9P 73.7 45 54 (0.8) 5,624 (0.5) 0.246
This workd (Acc.-DSP) {2, 3, 4} VU9P 70.9 40 12 (0.2) 7,637 (0.8) 0.130
@@
b On both Accuracy-DSP and Accuracy-LUT Pareto lines as shown in Fig. 21.
@@
e Dynamic power consumption reported by Vivado out-of-context (OOC) design flow. The static power is about 2.5W for all these designs.
@@ 5.9. Design-Space Exploration Strategies
Finding the optimal designs with grid search is time-consuming, as discussed in the previous section. This section investigates various DSE algorithms, including grid search, stochastic grid search (SGS), and Bayesian optimization. Fig. 20 presents the results. Each colored dot represents a design at a specific iteration, with different colors indicating different algorithms. The solid lines represent the Pareto Frontier of each algorithm over iterations, with each colored line corresponding to 22 iterations taking 40 hours. The uppermost grey line indicates a total of 343 iterations using extensive grid search over 624 hours, representing a baseline for comparison. The Bayesian optimization achieves similar results with just 22 iterations, significantly reducing processing time by a factor of 15.6 compared to Grid Search. Compared to SGS, Bayesian optimization’s efficient parameter search yields multiple points near the Pareto frontier, indicating its effectiveness in finding optimal designs and approaching global optima. This effective parameter search by employing past results demonstrates the robustness of the Bayesian optimization approach, enabling users to optimize models while minimizing time and effort.
5.10. Discussion
5.10.1. Comparative Evaluation.
@@ 5.10. Discussion
5.10.2. GNN: Scaling to graph-based models
Table 5 shows a case study on the GNN-based jet classifier of Odagiu et al. (Odagiu et al., 2024) with 8, 16, and 32 constituents on a Xilinx VU13P, where we apply the same PRUNING $O$-task with $\alpha_{p}=\beta_{p}=2\%$. MetaML-Pro automatically finds configurations that prune 53.1–65.7% of the parameters, reducing DSP usage by up to 2.5 times and LUT usage by up to about 2 times, while improving latency from 160–205 ns to 145 ns and the initiation interval from 15 to 10 ns, with accuracy degradations within roughly 1–2 percentage points relative to the baselines. These results extend our evaluation to a larger, irregular graph-based model family.
5.10.3. Additional Exploration Methods.
Beyond grid and Bayesian search, recent works explore learning-based and reinforcement learning (RL) strategies for accelerator DSE, as well as surrogate/ predictor-driven exploration. Examples include GNN- or RL-assisted HLS DSE and holistic AutoML-style co-design frameworks that learn fast proxies for hardware metrics to guide the search (Wu et al., 2021; Zhang et al., 2020; Reagen et al., 2017; Mehrabi et al., 2020; Kurek et al., 2016). In our setting, these methods target the same goal as BO-efficiently proposing promising design points under expensive evaluations, but differ in how they trade off sample efficiency, parallelism, and hyperparameter overhead. Our framework is compatible with these alternatives: the outer DSE loop interfaces our cross-stage O-tasks and K-task controls through the meta-model (CFG/LOG/MODEL-SPACE), so an evolutionary or RL agent can be substituted for BO without changing inner optimizations or the flow architecture.
Table 5. Performance comparison with FPGA designs of GNNs with 8/16/32 constituents on Xilinx VU13P FPGAs with a clock frequency of 200MHz. The pruning strategy is applied with both $\alpha_{p}$ and $\beta_{p}$ set to 2%.
Model FPGA Pru. (%) Acc.(%) Lat. (ns) II (ns) DSP (%) LUT (%) Pow. (W)a
GNN-8C (Odagiu et al., 2024) VU13P 0 64.9 160 15 2,191 (17.8) 472k (27.3) -
GNN-16C (Odagiu et al., 2024) VU13P 0 70.8 180 15 5,362 (43.6) 1,387k (80.3) -
GNN-32C (Odagiu et al., 2024) VU13P 50 75.8 205 15 2,120 (17.3) 1,162k (67.3) -
This work, GNN-8C VU13P 65.7 63.2 145 10 1,005 (8.2) 240k (13.9) 5.123
This work, GNN-16C VU13P 53.1 68.9 145 10 5,013 (40.8) 896k (51.9) 10.076
This work, GNN-32C VU13P 62.5 74.6 145 10 845 (6.9) 1,017k (58.8) 9.177
a On-chip power consumption estimated using Xilinx Vivado 2020.1 report_power in out-of-context (OOC) mode at 200 MHz, core logic only.
5.10.4. Extensibility of the Framework.
While this paper instantiates three SW-level O-tasks (SCALING, PRUNING) and an HLS-level mixed-precision QUANTIZATION (QHS), the pipe-task API and cross-stage controller are agnostic to the specific algorithm. The other optimization, such as hardware-aware NAS, knowledge distillation, structured sparsity and RTL level optimizations can be added as drop-in O-tasks/K-tasks that participate in the same FORK/BRANCH/REDUCE and BO-driven outer loop without changing the framework. In addition to new O/K-tasks, the outer DSE module is swappable: evolutionary or RL-based search can be plugged in as the controller that proposes $(\tau,\Pi)$, orchestrating the same FORK/BRANCH/REDUCE flow and inner searches without changing the pipeline code.
@@ 6. Conclusion
This paper presents a novel co-optimization framework for FPGA-based DNN accelerators, which comprises building blocks that facilitate rapid development of customized cross-stage design flows, automating the entire design iteration process. The results demonstrate that our approach significantly reduces DSP resource usage by up to 92% and LUT usage by up to 89%, while maintaining accuracy and without requiring human effort or domain expertise. In addition, our results reveal that Bayesian optimization in DSE significantly streamlines the process, achieving results comparable to grid search but with a 15.6-fold reduction in processing time. Future work will extend this approach to include RTL space and ASIC platforms, support a broader range of DNN architectures, including Variational Autoencoder (VAE) (Que et al., 2024c) and large language models (LLMs) (De Freitas et al., 2025), and incorporate more optimization strategies like balancing initiation interval (Que et al., 2021b; Rognlien et al., 2022) and distributed arithmetic (DA) (Sun et al., 2025b) for further improvements in hardware efficiency and model performance.
Acknowledgement: The support of the United Kingdom EPSRC (grant number EP/V028251/1, EP/N031768/1, EP/S030069/1, and EP/X036006/1), UKRI256, KIAT, Intel, and AMD is gratefully acknowledged. We thank Markus Rognlien, Filip Wojcicki and Shuo Liu and Anyan Zhao for their help.